OpenStudy (anonymous):

Find the standard equation of the plane containing the three points. Show All Work. (-2,1,1),(0,2,3),(1,0,-1)

OpenStudy (anonymous):

find two vector from that three points

OpenStudy (anonymous):

How do you do that?

OpenStudy (anonymous):

I will find one for you choose any two points (-2,1,1) -(0,2,3) <-2,-1,-2>

OpenStudy (anonymous):

Say you have... (2, -2, 1), (1, 0, -2), and (1, 1, 0) then your solution would be... The equation of a plane is Ax + By + Cz = D, where the normal vector to the plane is <A, B, C> to find the normal vector, you need two vectors that are in the plane. Taking the cross product of these two gives you the normal vector to the plane. To find two vectors in the plane, subtract two of these points from the other. So: (2, -2, 1) - (1, 0, -2) = <1, -2, 3> (1, 1, 0) - (1, 0, -2) = <0, 1, 2> Now take the cross product of these two vectors. You should get <-7, -2, 1> So your equation for the plane is -7x + -2y + z = D to solve for D, plug in one of the three points you're given -7(2) + -2(-2) + 1(1) = D so D = -9 Final equation: -7x - 2y + z = -9 You can multiply everything by a -1 if you want the x coefficient to be 7.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

How do you find the cross product?

OpenStudy (inkyvoyd):

(-2,1,1),(0,2,3),(1,0,-1)

OpenStudy (inkyvoyd):

There's an easier way.

OpenStudy (inkyvoyd):

Wel, not easier, just doesn't confuse you as much.

OpenStudy (anonymous):

Please explain it!

OpenStudy (inkyvoyd):

Do you know how to solve a system of linear equations with three unknowns?

OpenStudy (anonymous):

nope

OpenStudy (inkyvoyd):

Like ax+by+cz=0 dx+ey+fz=0 gx+hy+iz=0

OpenStudy (inkyvoyd):

where a,b,c,d,e,f,g,h, and i are known.

OpenStudy (anonymous):

wait I only need to find one equation

OpenStudy (anonymous):

wow um idk

OpenStudy (inkyvoyd):

nvm